OpenStudy (anonymous):

In a right triangle, the acute angles measure x+15 and 2x degrees. What is the measure of the larger acute angle of the triangle? 15 25 40 50

OpenStudy (welshfella):

the 2 acute angles add up to 90 degrees ( as the third angle = 90 degrees) so we have 2x + x +15 = 90 solve this equation for x then you can easily see which is the larger angle

OpenStudy (welshfella):

can you solve that equation?

OpenStudy (anonymous):

whats x ? @welshfella

OpenStudy (welshfella):

solve the equation to find x:- 2x + x + 15 = 90 3x + 15 = 90 now subtract 15 to both sides and simplify

OpenStudy (anonymous):

its 25. Thank you so much for the help ! @welshfella

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

x = 25 but the larger angle = 2*25 = 50